39. Use the medication order below to answer the next few questions GENERAL HOSPITAL PRESS HARD WITH BALLPOINT PEN. WRITE DATE&TIME AND SIGN EACH ORDER 316413 1/RS/O1 151 Ges Stree 0.3 ORDERS NOTED 7.30 FILLED BY DATE SIGNATURE M.D. PHYSICIANS ORDERS a.Examine the second medication order and calculate the flow rate for the infusion ordered. The drop factor is 10 drops per milliliter Examine the third medication order and state how you would administer the medication if the drop factor is 15 drops per milliter to infuse in two hours? 40. Use the label below to answer the following questions How many milliliters will you add to the bottle if you want 1,000,000 units per millliter? NDC 0019-0130-28n Rdter Pfizerpent penicillin G polassium) 3ans For Injection TWENTY MILLION UNITS Add 5,000,000 units of penicillin to 50 milliliters of 5% D/W; infuse in 30 minutes. The drop factor is 15 drops per milliliter. Calculate the flow rate in gtts/min. FORINTRAVENOU TT NFUSION ONL 7488
Answer #1

39) A. Formula to calculate the flow rate = total volume/ total time in minutes × drop factor

Total volume to be infused =750ml ,total time = 8hrs = 8×60=480mins.

Drop factor =10 drops /ml

Now, applying to the formula:-

= 750/ 480 × 10

= 15.6 gtt / min

39) B. Total volume to be infused =250ml

Drop factor = 15drops /ml

Total time to infuse is 2hours = 2×60= 120mins.

Flow rate = total volume /total time in mins × drop factor

= 250/ 120 ×15

= 31.25 gtt / min

40) A. 20 million units of pfizerpen is present in 100ml solution then , 1 million units of pfizerpen will be present in 100 / 20000000 = 0.000005 ml will be added to the bottle if you want 1000000 units per ml.

40.B) Total volume to be infused = 50 ml

Total time to infuse = 30 mins

Drop factor is 15 drops / ml

Formula to calculate flow rate = 50 /30 ×15 =25 gtt/ min